a) What current is needed to generate the magnetic field strength of 5.0*10^-5 T at a point 2.3 cm from a long, straight wire?

b) What current is needed to generate the magnetic field strength of 5.0*10^-3 T at a point 2.3 cm from a long, straight wire?

c) What current is needed to generate the magnetic field strength of 1 T at a point 2.3 cm from a long, straight wire?

d) What current is needed to generate the magnetic field strength of 10 T at a point 2.3 cm from a long, straight wire?
Answer #1
Concepts and reason

Concepts here will revolve around magnetic field at a point from a straight line, varying magnetic fields, and current associated with it.

Relation between current and magnetic is obtained from Biot-Savart law, which can be varied from case to case to get results.

Fundamentals

The magnetic field at a point at a distance from current carrying conductor is given by

B=μ0I2πdB = \frac{{{\mu _0}I}}{{2\pi d}}

Here BBis the magnetic field, μ0(=4π×107Hm){\mu _0}\left( { = 4\pi \times {{10}^{ - 7}}\;\frac{{\rm{H}}}{{\rm{m}}}} \right) is the permeability of free space, IIis the associated current, π(=3.14)\pi \left( { = 3.14} \right)is a constant and ddis the distance of the point from the rod.

(a)

Rearrange the expression B=μ0I2πdB = \frac{{{\mu _0}I}}{{2\pi d}} for current (I),

I=2πBdμ0I = \frac{{2\pi Bd}}{{{\mu _0}}}

The value of current due to the long straight wire can be calculated as,

I=2πBdμ0I = \frac{{2\pi Bd}}{{{\mu _0}}}

Convert the unit of distance from cm to m{\rm{m}}.

d=2.3cm=2.3cm(1m100cm)=0.023m\begin{array}{c}\\d = {\rm{2}}{\rm{.3}}\;{\rm{cm}}\\\\{\rm{ = 2}}{\rm{.3}}\,{\rm{cm}}\left( {\frac{{1{\rm{ m}}}}{{{\rm{100 cm}}}}} \right)\\\\ = 0.023\;{\rm{m}}\\\end{array}

Substitute 0.023m{\rm{0}}{\rm{.023}}\;{\rm{m}}for dd,5×105T{\rm{5 \times 1}}{{\rm{0}}^{{\rm{ - 5}}}} {\rm{T}} for BB,4π×107TmA1{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}} for μ0{\mu _0} in the expression of magnetic field due to straight wire.

I=2π(5×105T)(0.023m)(4π×107TmA1)=5.75A\begin{array}{c}\\I = \frac{{{\rm{2\pi }}\left( {{\rm{5 \times 1}}{{\rm{0}}^{{\rm{ - 5}}}} {\rm{T}}} \right)\left( {{\rm{0}}{\rm{.023}} {\rm{m}}} \right)}}{{\left( {{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}}} \right)}}\\\\ = 5.75 {\rm{A}}\\\end{array}

(b)

The value of current due to the long straight wire can be calculated as,

I=2πBdμ0I = \frac{{2\pi Bd}}{{{\mu _0}}}

Convert the unit of distance cm{\rm{cm}}to m{\rm{m}}.

d=2.3cm=2.3cm(1m100cm)=0.023m\begin{array}{c}\\d = {\rm{2}}{\rm{.3}}\;{\rm{cm}}\\\\{\rm{ = 2}}{\rm{.3}}\,{\rm{cm}}\left( {\frac{{1{\rm{ m}}}}{{{\rm{100 cm}}}}} \right)\\\\ = 0.023\;{\rm{m}}\\\end{array}

Substitute 0.023m{\rm{0}}{\rm{.023}}\;{\rm{m}}for dd, 5×103T{\rm{5 \times 1}}{{\rm{0}}^{{\rm{ - 3}}}} {\rm{T}}for BB,4π×107TmA1{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}} for μ0{\mu _0} in the expression of magnetic field due to straight wire.

I=2π(5×103T)(0.023m)(4π×107TmA1)=575A\begin{array}{c}\\I = \frac{{{\rm{2\pi }}\left( {{\rm{5 \times 1}}{{\rm{0}}^{{\rm{ - 3}}}} {\rm{T}}} \right)\left( {{\rm{0}}{\rm{.023}} {\rm{m}}} \right)}}{{\left( {{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}}} \right)}}\\\\ = 575 {\rm{A}}\\\end{array}

(c)

The value of current due to the long straight wire can be calculated as,

I=2πBdμ0I = \frac{{2\pi Bd}}{{{\mu _0}}}

Convert the unit of distance from cm to m{\rm{m}}.

d=2.3cm=2.3cm(1m100cm)=0.023m\begin{array}{c}\\d = {\rm{2}}{\rm{.3}}\;{\rm{cm}}\\\\{\rm{ = 2}}{\rm{.3}}\,{\rm{cm}}\left( {\frac{{1{\rm{ m}}}}{{{\rm{100 cm}}}}} \right)\\\\ = 0.023\;{\rm{m}}\\\end{array}

Substitute 0.023m{\rm{0}}{\rm{.023}}\;{\rm{m}}for dd, 1T{\rm{1}} {\rm{T}}forBB,4π×107TmA1{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}} forμ0{\mu _0} in the expression of magnetic field due to straight wire.

I=2π(1T)(0.023m)(4π×107TmA1)=115×103A\begin{array}{c}\\I = \frac{{{\rm{2\pi }}\left( {{\rm{1}} {\rm{T}}} \right)\left( {{\rm{0}}{\rm{.023}} {\rm{m}}} \right)}}{{\left( {{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}}} \right)}}\\\\ = 115 \times {10^3} {\rm{A}}\\\end{array}

(d)

The value of current due to the long straight wire can be calculated as,

I=2πBdμ0I = \frac{{2\pi Bd}}{{{\mu _0}}}

Convert the unit of distance from cm to m{\rm{m}}.

d=2.3cm=2.3cm(1m100cm)=0.023m\begin{array}{c}\\d = {\rm{2}}{\rm{.3}}\;{\rm{cm}}\\\\{\rm{ = 2}}{\rm{.3}}\,{\rm{cm}}\left( {\frac{{1{\rm{ m}}}}{{{\rm{100 cm}}}}} \right)\\\\ = 0.023\;{\rm{m}}\\\end{array}

Substitute 0.023m{\rm{0}}{\rm{.023}}\;{\rm{m}}for dd,10T{\rm{10}} {\rm{T}}forBB,4π×107TmA1{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}} forμ0{\mu _0} in the expression of magnetic field due to straight wire.

I=2π(10T)(0.023m)(4π×107TmA1)=115×104A\begin{array}{c}\\I = \frac{{{\rm{2\pi }}\left( {{\rm{10}} {\rm{T}}} \right)\left( {{\rm{0}}{\rm{.023}} {\rm{m}}} \right)}}{{\left( {{\rm{4\pi \times 1}}{{\rm{0}}^{{\rm{ - 7}}}} {\rm{T}} \cdot {\rm{m}} \cdot {{\rm{A}}^{ - 1}}} \right)}}\\\\ = 115 \times {10^4} {\rm{A}}\\\end{array}

Ans: Part a

The value of current due to the long straight wire is 5.75A5.75{\rm{ A}}.

Part b

The value of current due to the long straight wire is 575A575{\rm{ A}}.

Part c

The value of current due to the long straight wire is 115×103A115 \times {10^3}\;{\rm{A}}.

Part d

The value of current due to the long straight wire is 115×104A115 \times {10^{4\;}}\;{\rm{A}}.